how many moles of helium gas occupy 6.2 L at 303 K and 1.15 atm?
Answer #1

Ideal gas law , PV=nRT

No. of mol, n=PV/RT

From given data , n= (1.15*6.2)/(0.082*303)=0.28 mol

Thank you